3x-4=0 isnt this 3x=4 then x=3/4? The solution says its 1 1/3 or 4/3 but that isnt right? Please help fast. Thanks!
Answer by Alan3354(69443)

3x-4=0 isnt this 3x=4 then x=3/4?
------------
3x-4=0 isnt this
3x=4 is right.
Then divide both sides by 3
x = 4/3